Suppose you bought stock for $1,000 and sold it for $1,800 five year later. You paid 28% capital gain tax. What was your after-tax rate of return?
Answer #1

The purchase price was $1,000 and the net proceeds from sales was equal to (1-28%)*1800 = 1296. Find the rate of return

1296 = 1000(1+r)^5

(1296/1000)^(1/5) - 1 = r

This gives the after-tax rate of return at 5.322%